OpenStudy (anonymous):

A ball is thrown directly downward with an initial speed of 8.00 m/s, from a height of 30.0 m. After what time interval does it strike the ground?

OpenStudy (shane_b):

Initial velocity (Vi) = 8m/s Acceleration due to gravity (a) = 9.8 m/s^2 Distance (d)= 30m First calculate the velocity at the time of impact: \[V_f^2=V_i^2 + 2ad=(8m/s)^2 + 2(9.8m/s^2)(30m)=652 m/s\]\[V_f=\sqrt{652m/s}=25.5m/s\] Now that you know the final velocity, you can calculate the time using another kinematic equation: \[d=\frac{V_i+V_f}{2}t\]\[30=\frac{(8m/s + 25.5m/s)}{2}t\]\[t=1.79s\]

OpenStudy (lgbasallote):

you can just use the fomula \[\huge x - x_0 = V_0 t + \frac 12 at^2\] where: x = final height x _ 0 = initial height V_ 0 = initial speed t= time a = acceleration in this case x would be 30 and x_0 would be 0 because obvioulsy you START at the origin so you start with 0 height and then it says the height from where the ball was thrown is 30 then, the v_0 would be 8 m/s because it says the initial speed is 8 m/s and the acceleration would be 9.8 m/s^2 because it is the acceleration due to gravity so substitute \[\huge 30 - 0 = (8)t + \frac 12 (9.8)t^2\] simplify that.. \[\huge \implies 30 = 8t + 4.9t^2\] subtract 30 from both sides.. \[\huge \implies 4.9t^2 + 8 t - 30 = 0\] use quadratic formula to get the value of t \[\LARGE t = \frac{-8 \pm \sqrt{(8)^2 - 4(4.9)(-30)}}{2(4.9)}\] \[\huge \implies t \approx 1.78921 \implies t \approx 1.79 s\] does that help?

OpenStudy (shane_b):

Pfft....I wanted to avoid solving a quadratic :P

OpenStudy (lgbasallote):

wimp :P lol

OpenStudy (shane_b):

Yes...yes I am !

OpenStudy (anonymous):

Yes the above solutions give the same answer so either of those would do.